OpenStudy (anonymous):

a store is having a sale on 200 dvd movies. some of the movies are being sold for $10 and some of them are being sold for $12. if the store made $2200 after selling all 200 dvds, how many $12 dvds were sold

OpenStudy (aripotta):

ok so you need to set up a system of equations. we'll assign the variable "a" for the $10 movies and "b" of the $12 movies a + b = 200 since there are 200 movies total 10a + 12b = 2200 since the store made $2200 after selling them all

OpenStudy (aripotta):

what do we need to do now?

OpenStudy (anonymous):

im writing it down one sec

OpenStudy (anonymous):

i have no idea what to do next

OpenStudy (aripotta):

let's take the first equation and solve for a so we can plug it into the second equation

OpenStudy (anonymous):

whats the second equation

OpenStudy (aripotta):

a + b = 200 is the first and 10a + 12b = 2200 is the second

OpenStudy (anonymous):

o okay

OpenStudy (aripotta):

so using the first equation, what does a equal?

OpenStudy (anonymous):

one sec im going to look back at my notes

OpenStudy (aripotta):

a + b = 200 you have to get a by itself on the left. so to do that, you need to subtract b from both sides. a = 200 - b

OpenStudy (aripotta):

now that we know what a equals, we can plug that into the second equation 10a + 12b = 2200 10(200 - b) + 12b = 2200

OpenStudy (aripotta):

now you just need to distribute and solve for b and that will be your answer

OpenStudy (anonymous):